OpenStudy (anonymous):

find the equation of the tangent line to the curve when x has the given value. f(x) = sqrt (x) ; x= 49

OpenStudy (danjs):

Can you find the derivative of y = x^(1/2)

OpenStudy (danjs):

The line will be y - f(49) = f ' (49)*(x - 49)

OpenStudy (triciaal):

the slope of the tangent line is the first derivative of the curve at that point

OpenStudy (danjs):

\[\frac{ d }{ dx }\sqrt{x} = \frac{ d }{ dx }x ^{1/2} = \frac{ 1 }{ 2 }x ^{1/2 - 1} = \frac{ 1 }{ 2 }x ^{-1/2} = \frac{ 1 }{ 2\sqrt{x} }\]

OpenStudy (danjs):

Just used the power rule for derivatives. \[\frac{ d }{ dx }x^n = n*x ^{n-1}\]

OpenStudy (danjs):

With the point x=49, y = f(49) = sqrt(49) = 7 point; (49,7) and slope m = f '(49) = 1/(2*7) = 1/14 Form Line from Point-Slope Form y - f(49) = f '(49) *(x-49)

OpenStudy (danjs):

u understand everything... @Ko-Ki

OpenStudy (anonymous):

i think so thx

OpenStudy (danjs):

here is the graph

OpenStudy (anonymous):

thanks